# Coupling Electromagnetism to Torsion: Black Holes and Spin-Charge Interactions

### v1

@@ V Exact black hole solutions in 3 dimensions
Besides the interesting exact solution found in 4 dimensions, the non-minimal coupling between torsion and electromagnetism provides interesting features in three dimensions as deformation of the famous BTZ black hole Banados:1992wn . Three-dimensional black holes are important theoretical laboratories because gravity in three dimensions lacks local propagating degrees of freedom, allowing for a clearer analysis of the role played by global charges and topological effects. In particular, the BTZ black hole is a remarkable solution of (2+1)-dimensional gravity with a negative cosmological constant, featuring horizons, thermodynamic properties, and close connections to quantum gravity and holography, despite the absence of curvature singularities or gravitational radiation. Moreover, black hole solutions with torsion have been previously explored in various three-dimensional gravity models, such as Poincaré gauge theories and Chern–Simons-like formulations Blagojevic:2013bu ; Mielke:1991nn . In these contexts, torsion can contribute non-trivially to conserved charges, modify the thermodynamics of the BTZ black hole, and influence the corresponding holographic dual field theories. In contrast to those models, where torsion typically plays the role of an effective cosmological constant, we will show below that in our model, torsion contributes instead as a charge-like parameter. Our model thus offers a new framework to further investigate the role of torsion in three-dimensional black hole physics.
@@
To perform the outline analysis in a systematic way, we add a negative cosmological term to the Lagrangian (26), namely $L_{\Lambda}=-2\Lambda$ to include the standard BTZ solution Banados:1992wn as a limiting case, when the non-minimal couplings are set to zero. This extra term in the Lagrangian only introduces a term $-g_{\mu\nu}\Lambda$ in the tetrad equation (27).
@@ V.1 Spherically symmetric exact charged BTZ-like solution
$T_{1}$ $=$ $f(r)\left(\pm T_{4}(r)+\frac{1}{r}\right)\,,\quad T_{2}=\pm\frac{1}{2}K_{1}f(r)\,,\quad T_{3}=T_{5}=0\,,\quad T_{7}=K_{1}\,,\quad$ (58)
$T_{4}$ $=$ ±$\pm\frac{f^{\prime}(r)}{2f(r)}\mp\frac{1}{r}+\frac{K_{2}}{f(r)}+\frac{\kappa_{s}\log r}{f(r)}\,,\quad T_{6}=\mp\frac{K_{1}}{2f(r)}\,,\quad T_{8}=\pm f(r)\left(r-T_{9}(r)\right)\,,$ (59)
@@
where $q_{e}$ is the electric charge. Finally, the metric solution is an extension of the non-rotating charged BTZ black hole in 3-dimensions Martinez:1999qi , with a new coupling related to the intrinsic spin mediated by the new interaction term $\kappa_{\rm s}k_{3}q_{e}$:
@@
Compared to previously known three-dimensional solutions with torsion, where torsion typically appears as an effective constant or as a modification of the cosmological term Mielke:2003xx ; Garcia:2003nm ; Blagojevic:2003wn , the new feature our solution exhibits is dynamically coupled torsion components that evolve with the radial coordinate. Moreover, it gives rise to an intrinsic spin-charge hair in three dimensions and then, the thermodynamic behaviour of the black hole may be modified with respect to the non-rotating charged BTZ solution.
@@ V.2 Slowly rotating charged BTZ-like solution
±$\pm\frac{1}{f_{A}(r)}\left(\left(T_{9,A}(r)-3r\right)N_{\phi,B}(r)-\frac{rf_{A}^{\prime}(r)A_{\phi,B}^{\prime}(r)+r^{2}q_{e}N_{\phi,B}^{\prime}(r)+2q_{e}T_{7,B}(r)}{q_{e}}\right)\,,$
@@
$H_{1}(r)$ $=$ ±$\pm\frac{K_{1}A_{\phi,B}^{\prime}(r)}{2rq_{e}}+\frac{T_{4,A}(r)}{2K_{1}q_{e}}\Big{[}r\left(f_{A}^{\prime}(r)A_{\phi,B}^{\prime}(r)+rq_{e}N_{\phi,B}^{\prime}(r)\right)+f_{A}(r)\left(rA_{\phi,B}^{\prime\prime}(r)-A_{\phi,B}^{\prime}(r)\right)\Big{]}$ (71)
@@
±$\pm\frac{K_{1}}{2q_{e}T_{4,A}(r)}\Big{[}\frac{K_{1}A_{\phi,B}^{\prime}(r)}{2r}-\left(A_{t,B}^{\prime}(r)+rA_{t,B}^{\prime\prime}(r)\right)\Big{]}\,.$
@@
$T_{2,B}$ $=$ ±$\pm\frac{1}{2}f_{A}(r)\left(K_{3}T_{9,A}(r)+T_{7,B}(r)-2K_{3}r\right)-\frac{q_{m}f_{A}(r){}^{2}}{2q_{e}}\left(T_{4,A}(r)\pm\frac{1}{r}\right)\pm\frac{1}{2}K_{1}f_{B}(r)$ (78)
@@
$T_{6,B}$ $=$ ±$\pm\frac{\left(K_{3}T_{9,A}(r)-T_{7,B}(r)\right)}{2f_{A}(r)}-\frac{K_{4}}{2q_{e}f_{A}(r)}\left(r^{2}T_{4,A}(r)\pm\left(T_{9,A}(r)-\frac{r^{2}f_{A}^{\prime}(r)}{f_{A}(r)}+r\right)\right)$ (80)
@@
$T_{8,B}$ $=$ ±$\pm f_{A}(r)\left(-T_{9,B}(r)-\frac{K_{1}q_{m}}{2q_{e}}\right)\pm f_{B}(r)\left(r-T_{9,A}(r)\right)\mp\frac{K_{1}K_{4}r^{2}}{2q_{e}}\,,$ (81)
@@
±$\pm\left(2\left(T_{9,A}-r\right)T_{4,B}-\frac{rf_{B}\left(k_{3}^{2}q_{e}^{2}+32k_{1}k_{2}q_{e}^{2}+32k_{2}\Lambda r^{2}\right)}{16k_{2}k_{3}q_{e}f_{A}{}^{2}}\right)-\frac{K_{1}^{2}f_{B}}{2f_{A}{}^{2}}$
±$\pm T_{4,A}\left(2T_{9,B}+\frac{K_{1}K_{4}r^{2}}{q_{e}f_{A}}+\frac{K_{1}q_{m}}{q_{e}}\right)\,,$
@@
±$\pm k_{3}^{2}q_{m}f_{A}\Big{]}+f_{A}^{\prime}\Big{[}\pm 16k_{2}K_{4}r^{3}\left(k_{3}q_{e}T_{4,A}+2\right)+8k_{2}k_{3}K_{4}rq_{e}\left(2T_{9,A}+r\right)\Big{]}$
@@
In the context of the AdS/CFT correspondence, BTZ black holes play a central role as gravity duals of thermal states in two-dimensional conformal field theories (CFTs) Brown:1986nw . The presence of torsion, particularly when it is dynamically sourced and couples non-minimally to the electromagnetic field, suggests that the dual theory is no longer a standard CFT but rather a deformation of it, possibly one with spin current sources or broken parity.
In our model, torsion components are activated by both charge and rotation, and they backreact on the metric. This is expected to modify the boundary stress tensor and the holographic Ward identities, as it happens in Klemm:2007yu ; Blagojevic:2013bu . In particular, the asymmetry between electric and magnetic sectors due to the coupling mediated by $k_{3}$ implies that the dual theory could distinguish between left- and right-movers, suggesting a chirally asymmetric dual CFT Castro:2014tta ; Detournay:2012pc . Moreover, the spin-charge hair sourced by torsion is expected to affect boundary transport phenomena associated with geometry, such as Hall viscosity or thermal Hall response. In particular, torsion has been shown to couple to spin currents and induce dissipationless transport coefficients in topological phases of matter Hughes:2011hv ; Hughes:2012vg ; Gromov:2014vla . These effects suggest that the holographic dual theory may exhibit torsional anomalies or contain new geometric response terms beyond the usual energy-momentum and electromagnetic current structure.
Finally, since torsion modifies the thermodynamics and potentially the central charges (via boundary charges in Chern–Simons-like formulations), the Cardy formula may receive corrections as in Park:2006gt ; Blagojevic:2006jk ; Blagojevic:2006hh . This could affect the microscopic counting of black hole entropy in the dual theory, an aspect worth exploring in future work.
@@ VI Conclusions
In order to find black hole solutions in 4-dimensions we assumed spherical symmetry for all involved fields. We found that in this case, the electromagnetic potential is the Coulomb potential, torsion is non-trivial, and the spacetime metric becomes a spin-charged generalisation of the Reissner–Nordström solution, which we presented in Eq. (51). The radial dependence of the metric in the new solution has the same form as that of the Reissner–Nordström solution, but with an effective electric charge. Novel features emerge from the total charge. In our new solution, this charge is not only given by the electric charge $q$, but by a combination of electric charge, the spin charge $\kappa_{s}$ and the coupling constants $k_{2}$ and $k_{3}$, which induce dynamical torsion and the non-minimal coupling between torsion and electromagnetism. The most intriguing feature of the solution is that, depending on the strength of the couplings and the size of the spin-charge, the total charge need not to be positive, which changes the phenomenology compared to the usual Reissner–Nordström solution. A future interesting study will be to investigate the phenomenology of these spin-charged black holes, for example the behaviour of their accretion discs, as extension of existing studies on spin-charged accretion discs around general relativistic black holes Gimeno-Soler:2025jnl ; Lahiri:2023mwj .
In 3-dimensions, we added a negative cosmological constant to the theory in order to find new torsion-based generalization of the BTZ black hole Banados:1992wn . As in 4-dimension, it turns out that the source term of the Maxwell equation, the covariant divergence of the anti-symmetric part of the Riemann-Cartan Ricci tensor vanishes and thus the Maxwell equations are unchanged. For the metric, we found an electromagnetically and spin charged non-rotating BTZ black hole solution (62), which is the 3-dimensional analogue to the previously discussed 4-dimensional solution. It serves as a background solution for our extension to a slowly rotating solution, (86). It exhibits a remarkable feature absent in the original BTZ solution: a spin-charge–angular momentum coupling induced by the non-minimal coupling constant $k_{3}$. This result highlights the possibility of richer torsional contributions in rotating geometries, motivating further exploration in various dimensions and potentially revealing new holographic properties.
@@
When a non-trivial background for the metric is considered, such as Reissner-Nordström or Kerr-Newmann spacetime geometry without torsion and electromagnetic field, perturbations can be studied too determine torsional quasinormal that may provide new signatures of torsion, potentially detectable through observations of gravitational waves emitted by binary black hole mergers. Further, superradiance generated by our solution might trigger new effects related to torsion Brito:2015oca .
Further future directions include relaxing the constraint $k_{4}=k_{5}=k_{6}=0$ to uncover new black hole geometries with enricher torsional structures. On the cosmological side, it would be compelling to study solutions that incorporate torsion-induced modifications to the metric and electromagnetic sources, particularly in scenarios involving magnetic fields. It is well known that magnetic fields play an important role in the early universe Grasso:2000wj , and since our theory naturally extends Einstein–Maxwell with torsion, it would be interesting to analyse the evolution of primordial magnetic fields within this framework, with the aim of identifying new effects of torsion in the early universe. Another potentially significant direction is to investigate whether our theory can induce spontaneous parity violation at the level of linear perturbations, as commonly occurs in theories with torsion or nonmetricity Hohmann:2022wrk ; Aoki:2023sum . This mechanism could offer a natural explanation for the cosmic birefringence reported in Minami:2020odp . Furthermore, extending the analysis of rotation to four-dimensional solutions could clarify whether the coupling between intrinsic spin charge and angular momentum—observed in the three-dimensional case—also arises in higher-dimensional rotating black holes. Such studies could uncover new thermodynamic and geometric features induced by torsion. We leave all of these studies for the future.

### v2

@@ V Exact black hole solutions in 3 dimensions
Besides the interesting exact solution found in 4 dimensions, the non-minimal coupling between torsion and electromagnetism provides interesting features in three dimensions as deformation of the famous BTZ black hole Banados et al. (1992). Three-dimensional black holes are important theoretical laboratories because gravity in three dimensions lacks local propagating degrees of freedom, allowing for a clearer analysis of the role played by global charges and topological effects. In particular, the BTZ black hole is a remarkable solution of (2+1)-dimensional gravity with a negative cosmological constant, featuring horizons, thermodynamic properties, and close connections to quantum gravity and holography, despite the absence of curvature singularities or gravitational radiation. Moreover, black hole solutions with torsion have been previously explored in various three-dimensional gravity models, such as Poincaré gauge theories and Chern–Simons-like formulations Blagojevic et al. (2013); Mielke and Baekler (1991). In these contexts, torsion can contribute non-trivially to conserved charges, modify the thermodynamics of the BTZ black hole, and influence the corresponding holographic dual field theories. In contrast to those models, where torsion typically plays the role of an effective cosmological constant, we will show below that in our model, torsion contributes instead as a charge-like parameter. Our model thus offers a new framework to further investigate the role of torsion in three-dimensional black hole physics.
@@
To perform the outline analysis in a systematic way, we add a negative cosmological term to the Lagrangian (26), namely $L_{\Lambda}=-2\Lambda$ to include the standard BTZ solution Banados et al. (1992) as a limiting case, when the non-minimal couplings are set to zero. This extra term in the Lagrangian only introduces a term $g_{\mu\nu}\Lambda$ in the tetrad equation (27).
@@ V.1 Spherically symmetric exact charged BTZ-like solution
$T_{1}$ $=$ $f(r)\left(\pm T_{4}(r)+\frac{1}{r}\right)\,,\quad T_{2}=\pm\frac{1}{2}K_{1}f(r)\,,\quad T_{3}=T_{5}=0\,,\quad T_{7}=K_{1}\,$$\quad$ (58)
$T_{4}$ $=$ $\pm\frac{f^{\prime}(r)}{2f(r)}\mp\frac{1}{r}+\frac{K_{2}}{f(r)}+\frac{\kappa_{s}\log r}{f(r)}\,,\quad T_{6}=\mp\frac{K_{1}}{2f(r)}\,,\quad T_{8}=\pm f(r)\left(r-T_{9}(r)\right)\,$, (59)
@@
where $q_{e}$ is the electric charge. Finally, the metric solution is an extension of the non-rotating charged BTZ black hole in 3-dimensions Martinez et al. (2000), with a new coupling related to the intrinsic spin mediated by the new interaction term $\kappa_{\rm s}k_{3}q_{e}$:
@@
Compared to previously known three-dimensional solutions with torsion, where torsion typically appears as an effective constant or as a modification of the cosmological term Mielke and Rincon Maggiolo (2003); Garcia et al. (2003); Blagojevic and Vasilic (2003), the new feature our solution exhibits is dynamically coupled torsion components that evolve with the radial coordinate. Moreover, it gives rise to an intrinsic spin-charge hair in three dimensions and then, the thermodynamic behaviour of the black hole may be modified with respect to the non-rotating charged BTZ solution.
@@ V.2 Slowly rotating charged BTZ-like solution
$\pm\frac{1}{f_{A}(r)}\left(\left(T_{9,A}(r)-3r\right)N_{\phi,B}(r)-\frac{rf_{A}^{\prime}(r)A_{\phi,B}^{\prime}(r)+r^{2}q_{e}N_{\phi,B}^{\prime}(r)+2q_{e}T_{7,B}(r)}{q_{e}}\right)\,$,
@@
$H_{1}(r)$ $=$ $\pm\frac{K_{1}A_{\phi,B}^{\prime}(r)}{2rq_{e}}+\frac{T_{4,A}(r)}{2K_{1}q_{e}}\Big[r\left(f_{A}^{\prime}(r)A_{\phi,B}^{\prime}(r)+rq_{e}N_{\phi,B}^{\prime}(r)\right)+f_{A}(r)\left(rA_{\phi,B}^{\prime\prime}(r)-A_{\phi,B}^{\prime}(r)\right)\Big]$ (71)
@@
$\pm\frac{K_{1}}{2q_{e}T_{4,A}(r)}\Big[\frac{K_{1}A_{\phi,B}^{\prime}(r)}{2r}-\left(A_{t,B}^{\prime}(r)+rA_{t,B}^{\prime\prime}(r)\right)\Big]\,.$
@@
$T_{2,B}$ $=$ $\pm\frac{1}{2}f_{A}(r)\left(K_{3}T_{9,A}(r)+T_{7,B}(r)-2K_{3}r\right)-\frac{q_{m}f_{A}(r){}^{2}}{2q_{e}}\left(T_{4,A}(r)\pm\frac{1}{r}\right)\pm\frac{1}{2}K_{1}f_{B}(r)$ (78)
@@
$T_{6,B}$ $=$ $\pm\frac{\left(K_{3}T_{9,A}(r)-T_{7,B}(r)\right)}{2f_{A}(r)}-\frac{K_{4}}{2q_{e}f_{A}(r)}\left(r^{2}T_{4,A}(r)\pm\left(T_{9,A}(r)-\frac{r^{2}f_{A}^{\prime}(r)}{f_{A}(r)}+r\right)\right)$ (80)
@@
$T_{8,B}$ $=$ $\pm f_{A}(r)\left(-T_{9,B}(r)-\frac{K_{1}q_{m}}{2q_{e}}\right)\pm f_{B}(r)\left(r-T_{9,A}(r)\right)\mp\frac{K_{1}K_{4}r^{2}}{2q_{e}}\,$, (81)
@@
$\pm\left(2\left(T_{9,A}-r\right)T_{4,B}-\frac{rf_{B}\left(k_{3}^{2}q_{e}^{2}+32k_{1}k_{2}q_{e}^{2}+32k_{2}\Lambda r^{2}\right)}{16k_{2}k_{3}q_{e}f_{A}{}^{2}}\right)-\frac{K_{1}^{2}f_{B}}{2f_{A}{}^{2}}$
$\pm T_{4,A}\left(2T_{9,B}+\frac{K_{1}K_{4}r^{2}}{q_{e}f_{A}}+\frac{K_{1}q_{m}}{q_{e}}\right)\,$,
@@
$\pm k_{3}^{2}q_{m}f_{A}\Big]+f_{A}^{\prime}\Big[\pm 16k_{2}K_{4}r^{3}\left(k_{3}q_{e}T_{4,A}+2\right)+8k_{2}k_{3}K_{4}rq_{e}\left(2T_{9,A}+r\right)\Big]$
@@
In the context of the AdS/CFT correspondence, BTZ black holes play a central role as gravity duals of thermal states in two-dimensional conformal field theories (CFTs) Brown and Henneaux (1986). The presence of torsion, particularly when it is dynamically sourced and couples non-minimally to the electromagnetic field, suggests that the dual theory is no longer a standard CFT but rather a deformation of it, possibly one with spin current sources or broken parity.
In our model, torsion components are activated by both charge and rotation, and they backreact on the metric. This is expected to modify the boundary stress tensor and the holographic Ward identities, as it happens in Klemm and Tagliabue (2008); Blagojevic et al. (2013). In particular, the asymmetry between electric and magnetic sectors due to the coupling mediated by $k_{3}$ implies that the dual theory could distinguish between left- and right-movers, suggesting a chirally asymmetric dual CFT Castro et al. (2014); Detournay et al. (2012). Moreover, the spin-charge hair sourced by torsion is expected to affect boundary transport phenomena associated with geometry, such as Hall viscosity or thermal Hall response. In particular, torsion has been shown to couple to spin currents and induce dissipationless transport coefficients in topological phases of matter Hughes et al. (2011, 2013); Gromov and Abanov (2015). These effects suggest that the holographic dual theory may exhibit torsional anomalies or contain new geometric response terms beyond the usual energy-momentum and electromagnetic current structure.
Finally, since torsion modifies the thermodynamics and potentially the central charges (via boundary charges in Chern–Simons-like formulations), the Cardy formula may receive corrections as in Park (2008); Blagojevic and Cvetkovic (2006b, a). This could affect the microscopic counting of black hole entropy in the dual theory, an aspect worth exploring in future work.
@@ VI Conclusions
In order to find black hole solutions in 4-dimensions we assumed spherical symmetry for all involved fields. We found that in this case, the electromagnetic potential is the Coulomb potential, torsion is non-trivial, and the spacetime metric becomes a spin-charged generalisation of the Reissner–Nordström solution, which we presented in Eq. (51). The radial dependence of the metric in the new solution has the same form as that of the Reissner–Nordström solution, but with an effective electric charge. Novel features emerge from the total charge. In our new solution, this charge is not only given by the electric charge $q$, but by a combination of electric charge, the spin charge $\kappa_{s}$ and the coupling constants $k_{2}$ and $k_{3}$, which induce dynamical torsion and the non-minimal coupling between torsion and electromagnetism. The most intriguing feature of the solution is that, depending on the strength of the couplings and the size of the spin-charge, the total charge need not to be positive, which changes the phenomenology compared to the usual Reissner–Nordström solution. A future interesting study will be to investigate the phenomenology of these spin-charged black holes, for example the behaviour of their accretion discs, as extension of existing studies on spin-charged accretion discs around general relativistic black holes Gimeno-Soler et al. (2025); Lahiri and Laemmerzahl (2023).
In 3-dimensions, we added a negative cosmological constant to the theory in order to find new torsion-based generalization of the BTZ black hole Banados et al. (1992). As in 4-dimension, it turns out that the source term of the Maxwell equation, the covariant divergence of the anti-symmetric part of the Riemann-Cartan Ricci tensor vanishes and thus the Maxwell equations are unchanged. For the metric, we found an electromagnetically and spin charged non-rotating BTZ black hole solution (62), which is the 3-dimensional analogue to the previously discussed 4-dimensional solution. It serves as a background solution for our extension to a slowly rotating solution, (86). It exhibits a remarkable feature absent in the original BTZ solution: a spin-charge–angular momentum coupling induced by the non-minimal coupling constant $k_{3}$. This result highlights the possibility of richer torsional contributions in rotating geometries, motivating further exploration in various dimensions and potentially revealing new holographic properties.
@@
When a non-trivial background for the metric is considered, such as Reissner-Nordström or Kerr-Newmann spacetime geometry without torsion and electromagnetic field, perturbations can be studied too determine torsional quasinormal that may provide new signatures of torsion, potentially detectable through observations of gravitational waves emitted by binary black hole mergers. Further, superradiance generated by our solution might trigger new effects related to torsion Brito et al. (2015).
Further future directions include relaxing the constraint $k_{4}=k_{5}=k_{6}=0$ to uncover new black hole geometries with enricher torsional structures. On the cosmological side, it would be compelling to study solutions that incorporate torsion-induced modifications to the metric and electromagnetic sources, particularly in scenarios involving magnetic fields. It is well known that magnetic fields play an important role in the early universe Grasso and Rubinstein (2001), and since our theory naturally extends Einstein–Maxwell with torsion, it would be interesting to analyse the evolution of primordial magnetic fields within this framework, with the aim of identifying new effects of torsion in the early universe. Another potentially significant direction is to investigate whether our theory can induce spontaneous parity violation at the level of linear perturbations, as commonly occurs in theories with torsion or nonmetricity Hohmann and Pfeifer (2022); Aoki et al. (2024). This mechanism could offer a natural explanation for the cosmic birefringence reported in Minami and Komatsu (2020). Furthermore, extending the analysis of rotation to four-dimensional solutions could clarify whether the coupling between intrinsic spin charge and angular momentum—observed in the three-dimensional case—also arises in higher-dimensional rotating black holes. Such studies could uncover new thermodynamic and geometric features induced by torsion. We leave all of these studies for the future.
Finally, it would be interesting to analyse if our theory (mainly the coupling controlled by $k_{3}$) can arise from a higher-dimensional Einstein-Cartan (or higher-derivative curvature) gravity including the topological invariants (Euler, Pontryagin/Chern, and Nieh–Yan) and then perform a Kaluza–Klein dimensional reduction. This kind of reduction is known to generate in four dimensions non-minimal electrodynamics with curvature–electromagnetism cross terms (e.g. $R\,F^{2}$, $R_{\mu\nu}F^{\mu\alpha}F^{\nu}{}_{\alpha}$, etc.), see for instance Mueller-Hoissen (1988); Salam and Strathdee (1982); Dereli and Ucoluk (1990); Cremmer et al. (1978). Exploring whether an Einstein–Cartan parent theory with such topological densities can induce couplings of the schematic form $\tilde{R}^{\mu\nu}F_{\mu\nu}$ is a natural continuation of our work and will be addressed elsewhere.